OpenStudy (anonymous):

just verifying \[\int tsec^{2}2tdt=\frac{t}{2}tan2t-\frac{1}{4}ln|sec2t| +C\]

OpenStudy (experimentx):

is this the integral??\[ \int t \; \sec^2 (2t) dt \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i did integration by parts

OpenStudy (experimentx):

Hmm .. let's try it by integration by parts. \[ \frac 12 \sec^2 (2t) t^2 - \int \int \sec^2 (2t) dt dt \\ \text{ Ignoring frst term}\\ \frac 12\int \tan (2t) dt = \frac 14 \ln |\sec (2t)|\] I guess it should be \[ \frac 12 \sec^2 (2t) t^2 - \frac 14 \ln |\sec (2t)|\]

OpenStudy (experimentx):

let's try asking wolf

OpenStudy (experimentx):

what went wrong http://www.wolframalpha.com/input/?i=int+t+sec^2+%282t%29

OpenStudy (anonymous):

I disagree with wolf

OpenStudy (experimentx):

oh ... i got the wrong formula for integration by parts. the first term should be \[ t \int \sec^2 (2t) = \frac 12t \tan (2t) \]

OpenStudy (experimentx):

why??

OpenStudy (experimentx):

it seem you are right ... it's just manipulation of log - log(sec x) = log(cos x)

OpenStudy (anonymous):

lets see u=t dv=sec^(2) 2t dt du=dt v=(1/2)tan2t \[uv-\int vdu\] \[\int tsec^{2}2tdt=\frac{t}{2}tan2t-\frac{1}{4}ln|sec2t| +C\]

OpenStudy (experimentx):

\[ - \frac 14 \ln |\sec (2t)| = \frac 14 \ln \left |\frac 1{\sec (2t)} \right | = \frac 14 \ln |\cos (2t)|\]

OpenStudy (anonymous):

oh....i see

OpenStudy (experimentx):

both answers are equivalent.

OpenStudy (anonymous):

makes sense.. thanks again experimentX!

OpenStudy (experimentx):

you are welcome :)